# Marginal and conditional summary measures: transportability and compatibility

### v1

@@ 3.2 Linear heterogeneous CTEX
The homogeneous illustrative model in Section 3.1 is sometimes considered too restrictive, in which case one can postulate an outcome model that accommodates for treatment effect heterogeneity across individuals. The heterogeneous illustrative model specifies the conditional expectation of potential outcome $Y^{t}$ for treatment $T=t$ given $X$ as:
@@
For a given link function, substituting Equation 6 into Equation 2 gives $\textnormal{CTEX}=\beta_{T}+\beta_{XT}x$ at $X=x$ on the linear predictor scale. That is, for the identity, logarithmic or logit link, the corresponding conditional mean difference, log risk ratio or log odds ratio, respectively, depends on the level of $X$. As such, $\beta_{T}$ is only the conditional treatment effect evaluated at $X=0$, and there is no longer a single CTEX estimand. For a given link function, substituting Equation 6 into Equation 3 yields $\textnormal{CTEM}=\beta_{T}+\beta_{XT}\bar{X}$. Similarly, substituting Equation 6 into Equation 4 gives $\textnormal{PACTE}=E_{X}\left(\beta_{T}+\beta_{XT}X\right)=\beta_{T}+\beta_{XT}E_{X}\left(X\right)=\beta_{T}+\beta_{XT}\bar{X}$.
@@
With respect to the marginal estimand, consider that $g(\cdot)$ is the identity link, such that the outcome-generating model is linear. On the mean difference scale, by the law of total expectation, $\textnormal{MTE}=\beta_{T}+\beta_{XT}\bar{X}$ (Remiro-Azócar, 2024). While $\beta_{T}$ no longer has a marginal interpretation, the marginal estimand coincides with the CTEM and PACTE, and can still be expressed in terms of model coefficients and the unconditional expectation of $X$. This is because the mean difference is directly collapsible (Huitfeldt et al., 2019). Namely, the marginal measure is equal to a weighted average of conditional measures, with the weights given by marginal distributions of the covariates that are conditioned on, e.g., for binary or categorical covariates, the weights would be the covariate proportions (Colnet et al., 2023).
@@
Unlike the mean difference, the (log) risk ratio is not directly collapsible (Kiefer & Mayer, 2019). The collapsibility weights for this summary measure are more complex; the marginal (log) risk ratio is not generally equal to a weighted average of conditional (log) risk ratios, if the weights are given by marginal covariate summary moments (Huitfeldt et al., 2019). Consequently, with $g(\cdot)=\log(\cdot)$ and the heterogeneous log-linear generative model, the expression for the marginal (log) risk ratio cannot simply be reduced to one involving $\beta_{T}$, $\beta_{XT}$ and $\bar{X}$. In this scenario, the MTE on the log risk ratio scale is not equivalent to the CTEM or the PACTE (Remiro-Azócar, 2024).
@@
Moreover, where there are multiple prognostic baseline covariates and in the presence of treatment effect heterogeneity, the marginal (log) risk ratio cannot be identified using the marginal distributions of the effect modifiers (Remiro-Azócar, 2024). Rather, one requires the full joint (multivariate) distribution of both effect modifiers and purely prognostic covariates that are associated with the former. For non-collapsible measures such as the (log) odds ratio, the MTE depends on the full joint distribution of purely prognostic covariates, even if these are not associated with any effect modifiers or in the absence of treatment effect heterogeneity.
@@
Table 3 presents a taxonomy of marginal estimands based on the heterogeneous illustrative models. Figure 2 indicates whether different estimands match for the heterogeneous illustrative models, across link functions and summary measures. The shaded squares denote matching estimand values; the dots denote the diagonal, where logically the estimands are equivalent. In Table 3 and Figure 2, the summary measures are on the additive linear predictor scale imposed by the link function used to model the corresponding outcome type.
@@
Figure 2: Matrices indicating whether different estimands are equivalent for the heterogeneous illustrative models. The shaded squares denote matching estimand values; the dots denote the diagonal, where estimands are equivalent by definition.
@@ 3.3 Non-linear (quadratic) heterogeneous CTEX
In the heterogeneous illustrative model in Section 3.2, the conditional outcome expectation on the linear predictor scale varies linearly with $X$, in both the active treatment and the control groups. Hence, the CTEX – the difference between the (transformed) conditional outcome expectations for each treatment – is also a linear function of $X$. In practice, such outcome expectations may vary non-linearly with the covariate in at least one of the treatment groups, in which case the CTEX is a non-linear function of $X$. We will now adopt another heterogeneous illustrative model, where the conditional outcome expectations on the linear predictor scale vary quadratically with $X$ in both treatment groups, such that the CTEX is a quadratic function of $X$.
@@
Again, $X$ is prognostic of outcome and a (conditional) treatment effect modifier on the linear predictor scale. In this case, the effect modification mechanism is more complex. Due to the second-order interaction – that is, the squared covariate-treatment product term – subjects with extreme values of $X$ experience lesser treatment efficacy, whereas subjects with moderate covariate values experience greater treatment efficacy, or vice versa (Mayer et al., 2017).
For a given link function, substituting Equation 7 into Equation 2 yields $\textnormal{CTEX}=\beta_{T}+\beta_{1T}x+\beta_{2T}x^{2}$ at $X=x$ on the linear predictor scale. That is, for the identity, logarithmic or logit link, the corresponding conditional mean difference, log risk ratio or log odds ratio, respectively, depends on the values of the covariate and the squared covariate. For a given link function, applying Equation 7 to Equation 3 yields $\textnormal{CTEM}=\beta_{T}+\beta_{1T}\bar{X}+\beta_{2T}\bar{X}^{2}$, a quadratic function of the covariate mean.
[…]
What we seek to show with the quadratic illustrative model is that, unlike for the model in Section 3.2, the CTEM and the PACTE are not generally equivalent. Where the CTEX is constant or a linear function of $X$, it just so happens that the PACTE expression collapses to the CTEM on the linear predictor scale. Nevertheless, this is not the case where the CTEX is a non-linear function of $X$. Here, the interpretation of the CTEM is restricted to the covariate mean and no longer reflects a population-average conditional effect. Likewise, and contrary to suggestions in the literature (Remiro-Azócar et al., 2021a), the PACTE is a population-average estimand that does not necessarily represent the effect for subjects with the mean covariate values.
With respect to the marginal estimand, consider that $g(\cdot)$ is the identity link. On the mean difference scale, by the law of total expectation:
@@
such that the marginal treatment effect estimand coincides with the PACTE but not with the CTEM. Moreover, making use of the formula $Var(X)=E(X^{2})-\bar{X}^{2}$, yields the marginal mean difference $\textnormal{MTE}=\beta_{T}+\beta_{1T}\bar{X}+\beta_{2T}\left(\bar{X}^{2}+Var(X)\right)$. Consequently, and due to the direct collapsibility of the mean difference, the marginal estimand can be reduced to simple terms involving model coefficients and the mean and variance of the effect modifier $X$.
Conversely, as per Section 3.2, this is not the case for the marginal (log) risk ratio with $g(\cdot)=\log(\cdot)$, and for the marginal (log) odds ratio with $g(\cdot)=\operatorname{logit}(\cdot)$, for which the MTE on the corresponding linear predictor scale is not equivalent to the CTEM or to the PACTE. Again, for these outcome-generating models and summary measures, the MTE depends on the full distribution of covariate $X$ (not just its mean and its variance), and also the distribution of baseline risk, as prognostic factors may not even be known or measured. With multiple prognostic baseline covariates and in the presence of treatment effect heterogeneity, the MTE will generally depend on the full joint distribution of both effect modifiers and purely prognostic covariates (Remiro-Azócar, 2024).
Table 4 presents a taxonomy of marginal estimands based on the quadratic illustrative models. Figure 3 indicates whether different estimands match for the quadratic illustrative models, across link functions and summary measures. The shaded squares denote matching estimand values; the dots denote the diagonal, where logically the estimands are equivalent. In Table 4 and Figure 3, the summary measures are on the additive linear predictor scale imposed by the link function used to model the corresponding outcome type.
@@
Figure 3: Matrices indicating whether different estimands are equivalent for the quadratic (heterogeneous) illustrative models. The shaded squares denote matching estimand values; the dots denote the diagonal, where estimands are equivalent by definition.
4 The case of indirect treatment comparisons
We have clarified the interpretation of different marginal and conditional summary measures, considering different types of outcomes and outcome-generating mechanisms, with and without effect modification. In order to illustrate potential implications of estimand incompatibility for evidence synthesis, we consider the following two-study scenario, increasingly encountered in health technology assessment (Phillippo et al., 2016, 2018).
A manufacturer submitting evidence for reimbursement has conducted its own “index” RCT comparing the efficacy of novel active treatment $A$ versus control $C$. Another company has previously developed a competing active treatment and performed an RCT (the “competitor” trial) comparing the efficacy of its product, active treatment $B$, against control $C$. An indirect treatment comparison between $A$ and $B$, anchored relative to common comparator $C$, is required for reimbursement purposes. Typically, the manufacturer submitting the evidence has individual patient data (IPD) from the index RCT, but only aggregate-level summary data from the competitor RCT. For the latter trial, access to subject-level data is unavailable due to privacy and confidentiality concerns. This typically restricts the target population of the analysis to the competitor trial population.
@@ 4 The case of indirect treatment comparisons
To produce the estimate $\hat{\Delta}_{AC}^{(BC)}$ in Equation 8, several population adjustment methods have been developed: (1) matching-adjusted indirect comparison (MAIC) (Signorovitch et al., 2010); (2) simulated treatment comparison (STC) (Ishak et al., 2015); (3) multilevel network meta-regression (ML-NMR) (Phillippo et al., 2020); (4) network meta-interpolation (NMI) (Harari et al., 2023); and (5) cross-network meta-regression (cross-NMR) (Hamza et al., 2023, 2024). An oft-overlooked aspect is that these estimate different summary measures, as outlined in Table 5. MAIC estimates the MTE; the “plug-in” or “mean-centering” STC version (Phillippo et al., 2016) estimates the CTEM, but simulation- or G-computation-based adaptations (Remiro-Azócar et al., 2022) estimate the MTE; ML-NMR can estimate the PACTE or the MTE (Phillippo et al., 2023); and NMI and cross-NMR estimate the CTEM.
@@
Matching-adjusted indirect comparison MTE
Simulated treatment comparison (plug-in) CTEM
Simulated treatment comparison (G-computation) MTE
Multilevel network meta-regression PACTE and MTE
Network meta-interpolation CTEM
Cross-network meta-regression CTEM
In MAIC and the G-computation version of STC, $\hat{\Delta}_{AC}^{(BC)}$ is an estimate of the MTE. As such, the $\hat{\Delta}_{BC}^{(BC)}$ plugged into Equation 8 must estimate a compatible marginal summary measure. This is not problematic; a point estimate of the MTE is typically reported in RCT publications. Alternatively, “crude” unadjusted point estimates of the MTE can be derived from published aggregate-level outcome data, e.g., for binary outcomes, marginal risk differences, log risk ratios or log odds ratios can be calculated from event counts reported in two-by-two contingency tables.
The plug-in version of STC, NMI and cross-NMR plug in mean covariate values for the competitor trial population when estimating the covariate-adjusted treatment effect for $A$ versus $C$, such that $\hat{\Delta}_{AC}^{(BC)}$ is an estimate of the CTEM. However, any estimate $\hat{\Delta}_{BC}^{(BC)}$ that is available for $B$ versus $C$ is likely an incompatible marginal or conditional summary measure. RCTs often report estimates of conditional estimands, e.g., where the published effect estimate is the treatment coefficient of a multivariable regression of outcome on treatment and baseline covariates. For simplicity, our exposition in Section 2 and Section 3 has considered a single baseline covariate, and that conditional summary measures conditioning on said covariate. Nevertheless, conditioning on different sets of covariates leads to different conditional summary measures or estimands.
Trials with conditional estimates are likely to report effects conditioned on different covariate sets. These are not necessarily compatible across trials – e.g., for non-collapsible measures – even if the trials have identical target populations (Daniel et al., 2021). For a given target population, there are a multitude of PACTEs and CTEMs, one for every combination of baseline covariates considered for conditioning. For the plug-in STC and NMI, if conditional estimates are reported for the competitor trial, both $\hat{\Delta}_{AC}^{(BC)}$ and $\hat{\Delta}_{BC}^{(BC)}$ would have to condition on the same covariate set to guarantee estimand compatibility.
@@
Importantly, population-level decision-making requires population-average estimates that are relevant to the decision target population. The target population need not be represented by that of the competitor trial or may not be reflected by any of the studies in the evidence base. An advantage of ML-NMR is that it is not necessarily limited to produce treatment effect estimates in the competitor trial population, and can produce population-average marginal and conditional estimates in any target population of interest. This typically requires an additional identifying assumption to be made, which may be relaxed or assessed in larger evidence networks.(Phillippo et al., 2020, 2023).

### v3

@@ 3.2 Linear heterogeneous CTEX
The homogeneous illustrative model in Section 3.1 is sometimes too restrictive, in which case one can postulate an outcome model that accommodates for treatment effect heterogeneity across individuals. The heterogeneous illustrative model specifies the conditional expectation of potential outcome $Y^{t}$ for treatment $T=t$ given $X$ as:
@@
Substituting Equation 6 into Equation 2 gives $\textnormal{CTEX}=\beta_{T}+\beta_{XT}x$ at $X=x$ on the linear predictor scale. That is, for the identity, logarithmic or logit link, the corresponding conditional mean difference, log risk ratio or log odds ratio, respectively, depends on the level of $X$. As such, $\beta_{T}$ is only the conditional treatment effect evaluated at $X=0$, and there is no longer a single CTEX estimand. Substituting Equation 6 into Equation 3 yields $\textnormal{CTEM}=\beta_{T}+\beta_{XT}\bar{X}$. Similarly, substituting Equation 6 into Equation 4 gives $\textnormal{PACTE}=E_{X}\left(\beta_{T}+\beta_{XT}X\right)=\beta_{T}+\beta_{XT}E_{X}\left(X\right)=\beta_{T}+\beta_{XT}\bar{X}$.
@@
With respect to the marginal estimand, suppose $g(\cdot)$ is the identity link, such that the outcome-generating model is linear. On the mean difference scale, by the law of total expectation, $\textnormal{MTE}=\beta_{T}+\beta_{XT}\bar{X}$. While $\beta_{T}$ no longer has a marginal interpretation, the marginal estimand coincides with the CTEM and PACTE, and can still be expressed in terms of model coefficients and the unconditional expectation of $X$. This is because the mean difference is directly collapsible (Huitfeldt et al., 2019). Namely, the marginal measure is equal to a weighted average of conditional measures, with the weights given by marginal distributions of the covariates that are conditioned on, e.g., for binary or categorical covariates, the weights would be the covariate proportions (Colnet et al., 2023).
@@
Unlike the mean difference, the (log) risk ratio is not directly collapsible (Kiefer & Mayer, 2019). The collapsibility weights for this summary measure are more complex; the marginal (log) risk ratio is in general not equal to a weighted average of conditional (log) risk ratios, if the weights are given by marginal covariate summary moments (Huitfeldt et al., 2019). Consequently, with $g(\cdot)=\log(\cdot)$ and the heterogeneous log-linear generative model, the marginal (log) risk ratio cannot be expressed in terms of $\beta_{T}$, $\beta_{XT}$ and $\bar{X}$. In this scenario, the MTE on the log risk ratio scale is not equivalent to the CTEM or the PACTE (Remiro-Azócar, 2024).
@@
Moreover, when there are multiple prognostic baseline covariates and in the presence of treatment effect heterogeneity, the marginal (log) risk ratio cannot be identified using the marginal distributions of the effect modifiers (Remiro-Azócar, 2024). Rather, one requires the full joint (multivariate) distribution of the effect modifiers and the purely prognostic covariates that are associated with the former. For non-collapsible measures such as the (log) odds ratio, the MTE depends on the full joint distribution of purely prognostic covariates, even if these are not associated with any effect modifiers or in the absence of treatment effect heterogeneity.
@@
Table 3 presents a taxonomy of marginal estimands based on the heterogeneous illustrative models. Figure 2 indicates whether different estimands match for the heterogeneous illustrative models, across link functions and summary measures. The blue squares denote matching estimand values; the dots denote the diagonal, where logically the estimands are equivalent. In Table 3 and Figure 2, the summary measures are on the additive linear predictor scale imposed by the link function used to model the corresponding outcome type.
@@
Figure 2: Matrices indicating whether different estimands are equivalent for the heterogeneous illustrative models. The blue squares denote matching estimand values; the dots denote the diagonal, where estimands are equivalent by definition.
@@ 3.3 Non-linear (quadratic) heterogeneous CTEX
In the heterogeneous illustrative model in Section 3.2, the conditional outcome expectation on the linear predictor scale varies linearly with $X$, in both the active treatment and the control groups. Hence, the CTEX – the difference between the (transformed) conditional outcome expectations for each treatment – is also a linear function of $X$. In practice, such outcome expectations may vary non-linearly with the covariate in at least one of the treatment groups, in which case the CTEX is a non-linear function of $X$. We now adopt another heterogeneous illustrative model, where the conditional outcome expectations on the linear predictor scale vary quadratically with $X$ in both treatment groups, such that the CTEX is a quadratic function of $X$.
@@
Again, $X$ is prognostic of outcome and a (conditional) treatment effect modifier on the linear predictor scale. In this case, the effect modification mechanism is more complex (Mayer et al., 2017). Due to the squared covariate-treatment product term, treatment efficacy depends on $X^{2}$. Substituting Equation 7 into Equation 2 yields $\textnormal{CTEX}=\beta_{T}+\beta_{1T}x+\beta_{2T}x^{2}$ at $X=x$ on the linear predictor scale. That is, for the identity, logarithmic or logit link, the corresponding conditional mean difference, log risk ratio or log odds ratio, respectively, depends on the values of the covariate and the squared covariate. Applying Equation 7 to Equation 3 yields $\textnormal{CTEM}=\beta_{T}+\beta_{1T}\bar{X}+\beta_{2T}\bar{X}^{2}$, a quadratic function of the covariate mean.
On the other hand, substituting Equation 7 into Equation 4 gives the population-average conditional estimand $\textnormal{PACTE}=E_{X}\left(\beta_{T}+\beta_{1T}X+\beta_{2T}X^{2}\right)=\beta_{T}+\beta_{1T}E\left(X\right)+\beta_{2T}E\left(X^{2}\right)=\beta_{T}+\beta_{1T}\bar{X}+\beta_{2T}E\left(X^{2}\right)$. The CTEM is no longer equal to the PACTE because the square of the covariate mean is not equal to the mean of the squared covariate. Because $f(X)=X^{2}$ is a convex function, due to Jensen’s inequality: $E(X^{2})\geq\bar{X}^{2}$. As both $E(X^{2})$ and $\bar{X}^{2}$ are positive, the inequality implies that $\textnormal{PACTE}>\textnormal{CTEM}$ for $\beta_{2T}>0$, and $\textnormal{PACTE}<\textnormal{CTEM}$ for $\beta_{2T}<0$.
With the quadratic illustrative model we want to show that, unlike for the model in Section 3.2, the CTEM and the PACTE are in general not equivalent. When the CTEX is constant or a linear function of $X$, the PACTE expression happens to collapse to the CTEM on the linear predictor scale. Nevertheless, this is not the case when the CTEX is a non-linear function of $X$. Here, the interpretation of the CTEM is restricted to the covariate mean and no longer reflects a population-average conditional effect. Likewise, and contrary to suggestions in the literature (Remiro-Azócar et al., 2021a), the PACTE is a population-average estimand that does not necessarily represent the effect for subjects with the mean covariate values.
With respect to the marginal estimand, suppose $g(\cdot)$ is the identity link. On the mean difference scale, by the law of total expectation:
@@
such that the marginal treatment effect estimand coincides with the PACTE but not with the CTEM. Moreover, the formula $Var(X)=E(X^{2})-\bar{X}^{2}$ implies that the marginal mean difference $\textnormal{MTE}=\beta_{T}+\beta_{1T}\bar{X}+\beta_{2T}\left(\bar{X}^{2}+Var(X)\right)$. Consequently, and due to the direct collapsibility of the mean difference, the marginal estimand can be reduced to simple terms involving model coefficients and the mean and variance of the effect modifier $X$.
Conversely, as per Section 3.2, this is not the case for the marginal (log) risk ratio with $g(\cdot)=\log(\cdot)$, and for the marginal (log) odds ratio with $g(\cdot)=\operatorname{logit}(\cdot)$ – nor for any other non-linear link function – for which the MTE on the corresponding linear predictor scale is not equivalent to the CTEM or to the PACTE. Again, for these outcome-generating models and summary measures, the MTE depends on the full distribution of covariate $X$ (not just its mean and its variance), and also the distribution of baseline risk, as some prognostic factors may be unavailable. With multiple prognostic baseline covariates and in the presence of treatment effect heterogeneity, the MTE will generally depend on the full joint distribution of the effect modifiers and purely prognostic covariates (Remiro-Azócar, 2024).
Table 4 presents a taxonomy of marginal estimands based on the quadratic illustrative models. Figure 3 indicates whether different estimands match for the quadratic illustrative models, across link functions and summary measures. The blue squares denote matching estimand values; the dots denote the diagonal, where logically the estimands are equivalent. In Table 4 and Figure 3, the summary measures are on the additive linear predictor scale imposed by the link function used to model the corresponding outcome type.
@@
Figure 3: Matrices indicating whether different estimands are equivalent for the quadratic (heterogeneous) illustrative models. The blue squares denote matching estimand values; the dots denote the diagonal, where estimands are equivalent by definition.
4 Indirect treatment comparisons
We have clarified the interpretation of different marginal and conditional summary measures, considering different types of outcomes and outcome-generating mechanisms, with and without effect modification. We illustrate potential implications of estimand incompatibility for evidence synthesis on the following two-study scenario, increasingly encountered in health technology assessment (Phillippo et al., 2016, 2018).
A manufacturer submitting evidence for reimbursement has conducted its own “index” RCT comparing the efficacy of novel active treatment $A$ versus control $C$. Another company developed earlier a competing active treatment and performed an RCT (the “competitor” trial) comparing the efficacy of its product, active treatment $B$, against control $C$. An indirect treatment comparison between $A$ and $B$, anchored at the common comparator $C$, is required for reimbursement purposes. Typically, the manufacturer submitting the evidence has individual patient data (IPD) from the index RCT, but only aggregate-level summary data from the competitor RCT. For the latter trial, access to subject-level data is unavailable due to privacy and confidentiality concerns, which restricts the target population of the analysis to the competitor trial population.
@@ 4 Indirect treatment comparisons
To produce the estimate $\hat{\Delta}_{AC}^{(BC)}$ in Equation 8, several population adjustment methods have been developed: matching-adjusted indirect comparison (MAIC) (Signorovitch et al., 2010), simulated treatment comparison (STC), multilevel network meta-regression (ML-NMR) (Phillippo et al., 2020), network meta-interpolation (NMI) (Harari et al., 2023) and cross-network meta-regression (cross-NMR) (Hamza et al., 2023, 2024). An oft-overlooked aspect is that these methods estimate different summary measures, as outlined in Table 5. The “plug-in” or “mean-centering” version of STC (STC-P) (Phillippo et al., 2016), NMI and cross-NMR estimate the CTEM; MAIC and simulation- or G-computation-based adaptations of STC (STC-G) (Remiro-Azócar et al., 2022) estimate the MTE; and ML-NMR can estimate the PACTE or the MTE (Phillippo et al., 2023).
@@
Matching-adjusted indirect comparison (MAIC) MTE
“Plug-in” simulated treatment comparison (STC-P) CTEM
“G-computation” simulated treatment comparison (STC-G) MTE
Multilevel network meta-regression (ML-NMR) PACTE and MTE
Network meta-interpolation (NMI) CTEM
Cross-network meta-regression (cross-NMR) CTEM
In MAIC and STC-G, $\hat{\Delta}_{AC}^{(BC)}$ is an estimate of the MTE. As such, the $\hat{\Delta}_{BC}^{(BC)}$ plugged into Equation 8 must estimate a compatible marginal summary measure. This is not problematic; a point estimate of the MTE is typically reported in RCT publications. Alternatively, “crude” unadjusted point estimates of the MTE can be derived from published aggregate-level outcome data, e.g., for binary outcomes, marginal risk differences, log risk ratios or log odds ratios can be calculated from event counts reported in $2\times 2$ contingency tables.
STC-P, NMI and cross-NMR substitute mean covariate values for the competitor trial population when estimating the covariate-adjusted treatment effect for $A$ versus $C$, such that $\hat{\Delta}_{AC}^{(BC)}$ is an estimate of the CTEM. However, any estimate $\hat{\Delta}_{BC}^{(BC)}$ that is available for $B$ versus $C$ is likely an incompatible marginal or conditional summary measure. RCTs often report estimates of conditional estimands, e.g., where the published effect estimate is the treatment coefficient of a multivariable regression of outcome on treatment and baseline covariates. For simplicity, our exposition in Section 2 and Section 3 considered a single baseline covariate, and all conditional summary measures condition on such covariate. However, conditioning on different sets of covariates leads to different conditional summary measures or estimands.
Trials with conditional estimates are likely to report effects conditioned on different covariate sets. These may be incompatible, especially for non-collapsible measures, even if the trials have identical target populations (Daniel et al., 2021). For a given target population, there are a multitude of PACTEs and CTEMs, one for every combination of baseline covariates used for conditioning. For STC-P, NMI and cross-NMR, if conditional estimates are reported for the competitor trial, $\hat{\Delta}_{AC}^{(BC)}$ and $\hat{\Delta}_{BC}^{(BC)}$ would have to condition on the same covariate set to guarantee estimand compatibility.
@@
Population-level decision-making requires population-average estimates that are relevant to the decision target population. The target population need not be represented by that of the competitor trial or may not be reflected by any of the studies in the evidence base. An advantage of ML-NMR is that it is not limited to produce treatment effect estimates in the competitor trial population, and can produce population-average marginal and conditional estimates in any target population of interest. This typically requires an additional identifying assumption, which may be relaxed or assessed in larger evidence networks.(Phillippo et al., 2020, 2023).